what is the solve for a+a+a=180 solve for a

x+(x+34)=90 solve for x

3(x-6)=2(x-9) solve for x

Answer by Paul(988)

a+a+a = 180
3a = 180
3a/3 = 180/3
a=60
x+(x+34) = 90
2x+34-34=90-34
2x=56
2x/2=56/2
x=28
3(x-6)=2(x-9)
3x-18 = 2x-18
3x-18-3x=2x-18-3x
-18=-18-x
-18+18=-18-x+18
0=x